Cranbrook Chemical Ltd. manufactures two industrial compounds. In the month of May, 15,000 litres of direct material costing $160,000 were processed at a cost of $400,000. The joint process yielded 16,000 containers of a compound known as Jarlon and 4,000 containers of a compound known as Kharton. The respective selling prices of Jarlon and Kharton are $38 and $58. Both products may be processed further. Jarlon may be processed into Jaxton at an incremental cost of $8 per jar of the final product while Kharton may be processed into Kraxton at an additional cost of $32 per jar of the final product. The volume of jars of the final product are: 12,000 and 3,000 for Jaxton and Kraxton respectively. The selling price of Jaxton is $48 per jar. The selling price of Kraxton is $102 per jar.


Using the sales value at splitoff method, the joint costs allocated to Jarlon would be

Select one:

A. $289,520.

B. $115,808.

C. $405,328.

D. $154,672.

E. $110,480.

Question 7

Advantages of the sales value at splitoff method include all of the following EXCEPT

Select one:

A. there is no anticipation of subsequent management decisions.

B. the allocation of joint costs could lead managers to make poor decisions.

C. it is simple.

D. it uses a meaningful denominator.

E. it does not presuppose an exact number of subsequent steps for further processing.

Question 8

Which of the following statements is TRUE regarding main products, byproducts, and scrap?

Select one:

A. Product classifications do not change over the short-run.

B. Product classifications do not change over the long-run.

C. The cause-and-effect criterion determines the classification.

D. The distinctions between main products, byproducts, and scrap are well-established in practice.

E. Product classifications may change over time.

Question 9

All of the following statements about the constant gross margin percentage of net realizable method are true EXCEPT

Select one:

A. some products may receive negative allocations of joint costs.

B. the gross margin percentage remains the same regardless of the different amounts of separable costs.

C. all products have equal gross margin percentages.

D. the gross margin is calculated by deducting all separable costs from revenue.

E. it is based on a tenuous underlying assumption.
Answer #1

1.Sales Value at split off :

Jarlon = 16,000*38 = $608,000

Kharton = 4,000*58 = $232,000

Joint Costs allocated to Jarlon = 400,000*608,000/840,000

= $289,524

i.e. A. $289,520 (approx.)

7. B. the allocation of joint costs could lead managers to make poor decisions.

It is not an advantage

8. E. Product classifications may change over time.

9.C. all products have equal gross margin percentages.

The gross margin percentages are not equal

Negative allocation is possible when gross margin is negative.